### v1

@@ Abstract
A longstanding open question in algorithm design is whether “combinatorial" matrix multiplication algorithms—avoiding Strassen-like divide-and-conquer—can achieve truly subcubic runtime $n^{3-\delta}$. We present an $O(n^{2.89})$-time exact algorithm, which only sums convolutions in $\mathbb{Z}_{m}^{k}$ (multivariate polynomial multiplications) via FFT, building on the work of Cohn, Kleinberg, Szegedy and Umans (CKSU’05). While the algorithm avoids recursion, the asymptotic speedup arises only for impractically large matrices.
Motivated by practical applications, we use this baseline to develop a new framework for fast approximate matrix multiplication (AMM), via low-degree approximations of the CKSU polynomials. We show that combining the aforementioned algorithm with black-box linear sketching already breaks the longstanding linear speed-accuracy tradeoff for AMM (Sarlos’06, Clarkson-Woodruff’13, Pagh’11, Cohn-Lewis’00), achieving $\frac{1}{r^{1.1}}\|\mathbf{A}\|_{F}^{2}\|\mathbf{B}\|_{F}^{2}$ error in $O(rn^{2})$-time.
Our main result is a low-degree approximation scheme for the CKSU polynomials, based on a Fourier-concentration lemma, yielding substantially smaller error in the distributional setting where $\mathbf{A},\mathbf{B}$ come from an i.i.d product-distribution; For random Gaussian matrices, this practical AMM algorithm attains smaller error than the best rank-$r$ SVD of the output matrix $\mathbf{A}\mathbf{B}$, in time $O(rn^{2})$. This is a substantial improvement over iterative Krylov subspace methods for low-rank approximation. Our theoretical and empirical results suggest the possibility of replacing MatMuls with sums of convolutions in LLM training and inference.
@@ 1 Introduction
The discovery of fast matrix multiplication (FMM) [Str86, Wil12, CW13, DWZ23]—asserting that the product $\mathbf{A}\mathbf{B}$ of two $n\times n$ real matrices can be computed in $n^{\omega}\sim O(n^{2.37})\ll n^{3}$ time [Str86, Wil12, Le ̵14, DWZ23]—had a profound impact on theoretical computer science and algorithm design. Alas, FMM algorithms are unlikely to be practical on any imaginable hardware, as they inherently rely on recursive divide-and-conquer schemes, which create memory and IO-bottlenecks [KAA20] and lead to highly asymptotic runtimes, earning them the infamous name “galactic algorithms" [LR13]. Another drawback of FMM algorithms in the context of data-driven applications, is that they are agnostic to the input matrices – recursion hinders the algorithm’s ability to exploit structure in the input matrices (e.g., sparsity or invariances), except in a black-box fashion [YZ05].
In attempt to capture practical algorithms, a concurrent line of work has focused on “combinatorial" MatMul algorithms [Arl+70, Wil07, Lin09, Cha15, Yu18, Abb+23], which avoid Strassen-like divide-and-conquer and algebraic decompositions, and have non-asymptotic runtimes [Bal+15, KS17]. While the term “combinatorial" is open to interpretation, it is ultimately meant to capture practical algorithms (more on this below). An outstanding open question in the field is whether combinatorial algorithms can achieve truly subcubic runtime, even for multiplying $n\times n$ boolean matrices :
Conjecture 1.1 ([Abb+23]).
@@
A recent breakthrough result by Abboud et al. [Abb+23] presented a combinatorial Boolean MatMul algorithm running in $O(n^{3}/2^{\Omega(\sqrt[7]{\log n})})$ time, highlighting the subtlety of the conjecture. However, for real matrices, the best known speedup using combinatorial techniques is much weaker, $O(n^{3}/\log^{2}n)$ [Wil07].
Bypassing 1.1 in pursuit of more practical algorithms, has initiated a long line of research on approximate matrix multiplication (AMM), which studies the best speed-accuracy tradeoff achievable by “combinatorial" algorithms in close to quadratic time. More precisely, for a prescribed parameter $r<n$, the goal is to produce, in $O(rn^{2})$ time, a matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$, which $\varepsilon$-approximates $\mathbf{A}\mathbf{B}$ in the Frobenius norm, where the (normalized) error $\varepsilon\rightarrow 0$ as $r\rightarrow n$. Essentially all known AMM algorithms use randomized sketching or sampling techniques [Sar06, DKM06, Pag13, MZ10, CW13, CNW16, CL99], and the state-of-art after more than 20 years of research is a linear speed-accuracy tradeoff:
@@
where $\mathbf{C}$ must be produced in $O(rn^{2})$ (randomized) time. We usually refer to the normalized error, given by $\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}$, where the above bound becomes $1/r$ (we call this the “linear bound”). The first error term can be obtained via the standard “sketch-and-solve" algorithm (e.g., CountSketch [CW13]), while the second one is obtained by a clever output-sensitive variation of CountSketch, using FFT (TensorSketch [Pag13]). The two bounds in are generally incomparable, but they coincide for the (hardest known) distribution of random Gaussian matrices. A conceptual limitation of all aforementioned AMM algorithms (except [CL99] which applies only to nonnegative matrices), is that they use compression techniques, i.e., produce a low-rank or sparse output matrix $\mathbf{C}$. In data-driven applications, most notably LLM training and inference, this is a severe limitation since compression crucially decreases the number of trainable parameters (see [Ail+25] and references therein). Moreover, a recent result of [OP24] proves that for compression-based algorithms, in a setup where Alice and Bob can send $O(rn)$ bits representing their respective input matrices $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times n}$ to a “referee" who must then compute the output based on their messages, the error bound in (AMM error) is tight for random Gaussian (or $\pm 1$) matrices.
Our work challenges 1.1, and sheds further light on the power of convolutions (which are omnipresent in practical algorithms) in the context of matrix-multiplication. We develop a new algorithmic framework for approximate matrix-multiplication, which relies only on several invocations of vanilla FFTs after small preprocessing, leading to improvements on the state-of-art in both exact and approximate regimes.
1.1 Our Results
Our work can be viewed as a variation of the group-theoretic framework of Cohn and Umans [CU03]— originally developped for fast matrix multiplication—to the approximate and combinatorial setting, i.e., restricted to abelian groups (polynomials), and disallowing recursive calls [^1] One could argue that FFT is itself a recursive algorithm, however, the divide-and-conquer used in FFT differs fundamentally from the recursion used in FMM algorithms – FFT is used only for the marginal encoding of $\mathbf{A}$ and $\mathbf{B}$ respectively, and not for the multiplication itself (as in FMM), see [Blä13]. Moreover, FFT is clearly used in practice, which is what “combinatorial” algorithms ultimately attempt to capture.: All the algorithms we present rely on fast polynomial multiplication, where our key departure from the Cohn-Umans setting is applying convolutions on low-degree approximations of the polynomials proposed by [Coh+05], while avoiding the use of the asymptotic sum inequality [Sch81]. While our work is merely an initial step toward practical MatMul  algorithms, we provide empirical evidence that our algorithm improves substantially on the best known speed-accuracy tradeoff of approximate matrix multiplication, making it suitable for data-driven applications such as inference speedup in deep neural networks [Ail+25].
Our first result is an exact “combinatorial" MatMul  algorithm, that runs in truly subcubic time:
Theorem 1.2 (Faster Combinatorial MM, Informal).
@@ 1.1 Our Results
The algorithm bootstraps a certain construction first proposed by [Coh+05] as a base tensor for FMM. In principle, the algorithm refutes 1.1, but it is unforuntately still highly asymptotic, beating naïve MatMul  ($\sim n^{3}$) only for very large matrices ($n\gtrsim 18^{12}$). This is our main motivation for focusing on the approximate convolutions, where the hope is to exploit the structure of the polynomials used in Theorem 1.2, in order to reduce the complexity of computing their convolution, possibly with some small accuracy loss.
As a proof-of-concept, we first show that a black-box combination of the exact combinatorial algorithm from Theorem 1.2 together with linear sketching, already yields the first super-linear speed-accuracy tradeoff for AMM, improving the state-of-art bound (AMM error):
@@
For any $r\leq n$, there exists a randomized algorithm that, given two square $n\times n$ matrices $\mathbf{A},\mathbf{B}$, produces in time $\tilde{O}(n^{2}r)$ a matrix $\mathbf{C}$, such that w.h.p, $\|\mathbf{C}-\mathbf{A}\mathbf{B}\|_{F}^{2}=O(\frac{1}{r^{1+\delta}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$ where $\delta\geq 0.1$.
We reiterate that FFT was already used in the AMM line of work (TensorSketch [Pag13]), albeit in a different way. Our work is the first to surpass the linear $1/r$ speed-accuracy tradeoff (AMM error) for AMM, without using FMM.
One suboptimal feature of Theorem 1.3 is that it is agnostic to the structure of the polynomials being convolved (the polynomials used in Theorem 1.2). The technical centerpiece of our work is a Fourier-concentration lemma for the CKSU polynomials, which leads to a non-blackbox AMM algorithm with substantially better tradeoff, for a large class of product distributions:
Theorem 1.4 (Fourier Truncation Algorithm, Informal version of Theorem 9.6, 9.13).
Let $\mathbf{A},\mathbf{B}$ be two independent $n\times n$ matrices whose entries are i.i.d. from sub-Gaussian distributions $\mathcal{D}_{A},\mathcal{D}_{B}$ with non-zero variance and second moments (e.g., Bernoulli, Rademacher, Gaussian). Then for any $r\leq n/2$, there exists a deterministic algorithm running in time $O(rn^{2})$ that computes a convolution of $3$-variate polynomials, outputting a matrix $\mathbf{C}$, such that $\mathbb{E}[\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}]\leq O\left(\frac{(n-2r)^{3}}{n^{4}}\right)$. In particular, for random $n\times n$ Gaussian or Rademacher matrices, the expected squared error is
$O\left(\left(\frac{n-2r}{n}\right)^{3}\right)\cdot\left\lVert\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}.$
The constants hidden in the $O(\cdot)$ notation depend on the distributions $\mathcal{D}_{A},\mathcal{D}_{B}$.
For calibration, notice that outputting the trivial expected solution $\mathbf{C}=\mathbb{E}[\mathbf{A}\mathbf{B}]=0^{n\times n}$ without even reading the inputs, already attains a squared error of $\left\lVert\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}$, and this is consistent with Theorem 1.4 when setting $r=0$. For Gaussian matrices, the error bound of Theorem 1.4 is smaller (by a factor of $\frac{(n-r)^{2}}{n^{2}}$) than the best rank-$r$ approximation of the output matrix $\mathbf{A}\mathbf{B}$, obtained by $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ [MP67] (see Figure Figure 5 for empirical validation). Note that even the fastest algorithms for approximating $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ to $\varepsilon$-additive accuracy require $\tilde{O}(rn^{2}/\sqrt{\varepsilon})$ time [MM15], via iterated matrix-vector (or block) multiplications ($\mathbf{A}\mathbf{B}x$), which makes these algorithms rather impractical. By contrast, the algorithm in Theorem 1.4 is independent of $\varepsilon$ or the spectral gap $(\sigma_{r+1}(\mathbf{A}\mathbf{B})/\sigma_{r}(\mathbf{A}\mathbf{B}))$, and is parallel (has depth $O(\log n)$ of FFT as opposed to depth $\tilde{O}(r/\sqrt{\varepsilon})$ of iterative Krylov-subspace algorithms). We present empirical evidence (see Figure 5) that corroborates this. We are not aware of non-iterative methods that can achieve $r$-$\mathsf{SVD}(\mathbf{A}\mathbf{B})$ even for random i.i.d matrices.
While the algorithm in Theorem 1.4 cannot be generalized to worst-case matrices (we give a simple counter-example, for which this method cannot work), finding the most general form of this method and adding randomness still remains a fascinating question with potential implications in practice.
A Lower Bound for Direct Convolution Algorithms
We complement our results by proving the limitations of using a single (exact) convolution for AMM. We prove that a single convolution over small abelian groups (e.g. low-degree polynomial multiplication) cannot surpass the linear speed-accuracy tradeoff in (AMM error). We also provide a simple algorithm which matches this error bound.
@@ A Lower Bound for Direct Convolution Algorithms
Theorem 1.5 generalizes the $\Omega(n^{3})$ lower bound of [CU03] on the size of abelian groups realizing exact matrix multiplication using convolutions, and provides an alternative, elementary algorithm achieving (AMM error), while possibly having better theoretical qualities such as high output rank.
2 Technical Overview
@@
Matrix Multiplication via Convolutions
@@
Unfortunately, the convolution algorithm in (2.1) is quite useless, as the polynomial $P_{\mathbf{A}}$ has degree $n^{3}$, hence the convolution (via FFT) requires $\Omega(n^{3}\log n)$ time, slower than naïve matrix multiplication(!). As such, [CU03] generalized the group-convolution approach (2.1) and the TPP property to general groups, where instead of polynomials, the matrices $\mathbf{A}$ and $\mathbf{B}$ are embedded into the “group algebra" $\mathbb{C}[G]$ of some general but fixed group $G$, and the two embeddings are then convolved inside $\mathbb{C}[G]$ to produce an element of $\mathbb{C}[G]$ containing the desired output $\mathbf{A}\mathbf{B}$. [^2] More formally, embed the indices $\mathbf{A}(i,j),\mathbf{B}(j^{\prime},k)\to G$ and compute $(\sum_{g\in G}\mathbf{A}_{g}g)\cdot(\sum_{h\in G}\mathbf{B}_{h}h)=\sum_{f\in G}(\sum_{gh=f}\mathbf{A}_{g}\mathbf{B}_{h})f$. See section 3 and [CU03] for more details. The TPP condition for general groups then becomes:
@@
The aforementioned $\Omega(n^{3})$ lower bound on the degree of TPP polynomials in (2.1) is not a coincidence – A simple proof [CU03] shows that any abelian group satisfying the TPP must have $\left|G\right|\geq n^{3}$, which seems to make abelian groups useless. In fact, our Theorem 1.5 generalizes this lower bound to the approximate setting, asserting that convolutions over abelian groups of size $|G|\leq rn^{2}$ cannot yield better than linear (normalized) Frobenius error $\Omega(1/r)$. The proof uses an elegant Cauchy-Schwarz argument on a carefully-chosen set of multiplicities (quadruplets). We show this is tight by providing a very simple and practical algorithm, using a single convolution in $\mathbb{Z}_{rn^{2}}$. The analysis is elementary, in contrast to sketching algorithms [Sar06, DKM06, CW13] which rely on the JL-moment property and Minkowski’s inequality to achieve (AMM error).
@@
The TPP lower bound for abelian groups shifted the focus of the group-theoretic-FMM line of work to finding non-abelian TPP groups of sub-cubic size [^3] In fact, an example of a group $G$ satisfying the TPP with $\left|G\right|=n^{2+o(1)}$ was given in [CU03]. [CU03, Coh+05]. Alas, non-abelian groups generally lack fast convolution algorithms, due to: (i) the absence of a fast Fourier transform algorithm; (ii) The fact that Fourier transform of non-abelian groups transforms convolutions into block-diagonal matrix products, as opposed to element-wise product in the abelian case (see section 3). This is where divide-and-conquer comes into play: applying the algorithm recursively on tiles (a-la Strassen) amortizes the cost of any linear transformation done before or after the recursive step, so applying the Fourier transform of $G$ is asymptotically free of charge. This is also where our approach departs from this line of work – Since we aim for non-recursive (“combinatorial”) algorithms, we must inevitably work with abelian groups which have FFTs. While this seems like a dead-end in light of the $\Omega(n^{3})$ degree lower bound for TPP polynomials, it turns out that multivariate polynomials (e.g., convolutions in $\mathbb{Z}_{m}^{k}$) satisfy a more general TPP condition, which allows to multiply block-matrices faster than the trivial time. We explain this idea, and how we use it to prove Theorem 1.2 next.
2.1 Exact Matrix Multiplication (Proof of Theorem 1.2)
Multivariate Polynomials and Simultaneous TPP
A followup work of Cohn, Kleinberg, Szegedy and Umans [Coh+05], proposed a simple generalization of the (TPP)., termed the Simultaneous Triple Product Property ( (STPP).), which captures the “capacity" of a group $G$ to compute $k$ independent matrix products using a single convolution. (Information-theoretically, this makes sense, since the TPP group-embedding e.g. (2.1) generates extremely sparse polynomials, so intuitively there’s plently of room to “pack" more information into the convolution, see Figure 1). Simply stated, the (STPP). describes when $G$ contains a collection of $d$ triples $(S_{i},T_{i},U_{i})_{i\in[d]}$ (for simplicity, assume $|S_{i}|=|T_{i}|=|U_{i}|=q$ are all equally sized), s.t (i) Each triplet marginally satisfies (TPP)., and (ii) The convolution in $G$ “doesn’t mix" the $d$ triplets ($(t_{j}-s_{j}^{\prime})+(u_{k}-t_{k}^{\prime})=(u_{i}^{\prime}-s_{i})=0$ implies $i=j=k\in[d]$, stated in additive notation). Hence, assuming $G$ has a fast Fourier transform, this property allows to compute $d$ independent $q\times q$ MatMuls for the price of a single convolution $\tilde{O}(|G|)$. Thus, if $|G|\ll dq^{3}$, this is faster than the naïve runtime. Unlike (TPP)., [Coh+05] show that there are nontrivial abelian groups satisfying (STPP)., and provide a simple non-trivial construction over the group $\mathbb{Z}_{m}^{3N}$ ($3N$-variate polynomials with bounded degree $m$ on each variable) with $d=2^{N}$ and $q=(m-1)^{N}$. This basic construction, which we’ll call “CKSU polynomials”, will play a central role in the proof of Theorem 1.4. In its most basic form, by setting $N=1$ (so d=2,q=(m1)d=2,q=(m-1)), it allows embedding two independent matrix products $\mathbf{A}_{1}\times\mathbf{B}_{1}$ and $\mathbf{A}_{2}\times\mathbf{B}_{2}$, of size $m-1\times m-1$ via the convolution of the following polynomials in $\mathbb{Z}_{m}^{3}$:

### v2

@@ Abstract
We study the capability of the Fast Fourier Transform (FFT) to accelerate exact and approximate matrix multiplication without using Strassen-like divide-and-conquer. We present a simple exact algorithm running in $O(n^{2.89})$ time, which only sums a few convolutions (FFTs) in $\mathbb{Z}_{m}^{k}$, building on the work of Cohn, Kleinberg, Szegedy and Umans [Coh+05]. As a corollary, combining this algorithm with linear sketching breaks the longstanding linear speed–accuracy tradeoff for “combinatorial” approximate matrix multiplication (AMM, [Pag13, Sar06, CW13]), achieving error $\tfrac{1}{r^{1.1}}\|\mathbf{A}\|_{F}^{2}\|\mathbf{B}\|_{F}^{2}$ in $O(rn^{2})$ time, using nothing but FFTs.
Motivated by the rich literature for approximating polynomials, our main contribution in this paper is extending the group-theoretic framework of Cohen and Umans (2003) to approximate matrix multiplication (AMM). Specifically, we introduce and study an approximate notion of the Triple Product Property, which in the abelian case is equivalent to finding a Sumset which minimizes (multi-)intersections with an arithmetic progression. We prove tight bounds on this quantity for abelian groups (yielding a simple and practical AMM algorithm via polynomial multiplication), and establish a weaker lower bound for non-abelian groups, extending a lemma of Gowers. Finally, we propose a concrete approach that uses low-degree approximation of multivariate polynomials for AMM, which we believe will lead to practical, non-asymptotic AMM algorithms in real-world applications, most notably LLM inference.
@@ 1 Introduction
The discovery of fast matrix multiplication (FMM) –asserting that the product $\mathbf{A}\mathbf{B}$ of two $n\times n$ real matrices can be computed in $n^{\omega}\sim O(n^{2.37})\ll n^{3}$ time [Str86, Wil12, Le ̵14, DWZ23]—had a profound impact on theoretical computer science and algorithm design. Alas, FMM algorithms are unlikely to be practical on any imaginable hardware, as they inherently rely on recursive divide-and-conquer schemes, which create memory and IO-bottlenecks [KAA20] and lead to highly asymptotic runtimes, earning them the infamous name “galactic algorithms" [LR13]. Another drawback of FMM algorithms in the context of data-driven applications, is that they are agnostic to the input matrices – recursion hinders the algorithm’s ability to exploit structure in the input matrices (e.g., sparsity or invariances), except in a black-box fashion [YZ05].
In attempt to capture practical algorithms, a concurrent line of work has focused on “combinatorial" MatMul algorithms [Arl+70, Wil07, Lin09, Cha15, Yu18, Abb+23], which avoid Strassen-like divide-and-conquer and algebraic decompositions, and have non-asymptotic runtimes [Bal+15, KS17]. While the term “combinatorial" is open to interpretation (especially when working over $\mathbb{R}$), we understand it as algorithms avoiding divide and conquer. An open question in the field is whether combinatorial algorithms can achieve truly subcubic runtime:
Conjecture 1.1 ([Abb+23, Sat94, Lee01]).
@@
Bypassing 1.1 in pursuit of more practical algorithms, has initiated a long line of research on approximate matrix multiplication (AMM), which studies the best speed-accuracy tradeoff achievable by “combinatorial” algorithms in close to quadratic time. More precisely, for a prescribed parameter $r<n$, the goal is to produce, in $O(rn^{2})$ time, a matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$, which $\varepsilon$-approximates $\mathbf{A}\mathbf{B}$ in the Frobenius norm, where the (normalized) error $\varepsilon\rightarrow 0$ as $r\rightarrow n$. Essentially all known AMM algorithms use randomized sketching or sampling techniques [Sar06, DKM06, Pag13, MZ10, CW13, CNW16, CL99], and the state-of-art after more than 20 years of research is a linear speed-accuracy tradeoff:
@@
where $\mathbf{C}$ must be produced in $O(rn^{2})$ (randomized) time. We usually refer to the normalized error, given by $\frac{\left\lVert\mathbf{C}-\mathbf{A}\mathbf{B}\right\rVert_{F}^{2}}{\left\lVert\mathbf{A}\right\rVert_{F}^{2}\cdot\left\lVert\mathbf{B}\right\rVert_{F}^{2}}$, where the above bound becomes $1/r$ (we call this the “linear bound”). The first error term can be obtained via the standard “sketch-and-solve” algorithm (e.g., CountSketch [CW13]), while the second one is obtained by a clever output-sensitive variation of CountSketch, using FFT (TensorSketch [Pag13]). The two bounds in (AMM error) are generally incomparable, but they coincide for the (hardest known) distribution of random Gaussian matrices. A conceptual limitation of all aforementioned AMM algorithms (except [CL99] which applies only to nonnegative matrices), is that they use compression techniques, i.e., produce a low-rank or sparse output matrix $\mathbf{C}$. In data-driven applications, most notably LLM training and inference, this is a severe limitation since compression crucially decreases the number of trainable parameters (see [Ail+25] and references therein). Moreover, a recent result of [OP24] proves that for compression-based algorithms, in a setup where Alice and Bob can send $O(rn)$ bits representing their respective input matrices $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times n}$ to a “referee” who must then compute the output based on their messages, the error bound in (AMM error) is tight for random Gaussian (or $\pm 1$) matrices.
1.1 Our Contributions
We view this work as an extension of the Group Theoretic approach developed by [CU03], originally for FMM, to the approximate setting. Put simply, we study AMM algorithms using (multi-variate) polynomial multiplication (and nothing else), avoiding Strassen style divide-and-conquer [^1] Divide and conquer in bilinear algorithms essentially boils down to the computation of a Kronecker power matrix with a vector, $\mathbf{U}^{\otimes k}v$, which can be done in $\Theta(mn^{k})$ assuming $\mathbf{U}$ is a $n\times m$ sized matrix with $m<n$, or $\Theta(kn^{k+1})$ when $m=n$. Suppose we treat $n$ and $m$ as constants, which leads the runtime to be asymptotically equivalent to the that of FFT (or even faster, when $m<n$, since $m$ does not scale with $k$). The key difference with FFT remains — FFT has widely available and highly optimized implementations with very small constants, and has non-asymptotic runtime. Thus, for real-life matrix sizes, it is incomparable in efficiency to generic bilinear algorithms..
Our first contribution is to put in writing a simple observation that allows one to obtain FFT-based algorithms from the group theoretic approach. Applying this observation on a construction proposed by [Coh+05] leads to the following result:
Theorem 1.2 (Faster “Combinatorial” MM).
@@ 1.1 Our Contributions
Unfortunately, despite FFT being practical, the algorithm is still highly asymptotic. This is our main motivation to explore approximate versions, where the hope is to exploit some structure of the underlying construction. As a proof-of-concept, we show that a black-box combination of the algorithm in Theorem 1.2 with standard linear sketching methods leads to the first super-linear speed-accuracy tradeoff for AMM:
@@
For any $r\leq n$, there exists a randomized algorithm that, given two square $n\times n$ matrices $\mathbf{A},\mathbf{B}$, produces in time $\tilde{O}(n^{2}r)$ a matrix $\mathbf{C}$, computed by using FFTs, such that w.h.p, $\|\mathbf{C}-\mathbf{A}\mathbf{B}\|_{F}^{2}=O(\frac{1}{r^{1+\delta}}\left\lVert\mathbf{A}\right\rVert_{F}^{2}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$ where $\delta\geq 0.1$.
We reiterate that FFT was already used in the AMM line of work (TensorSketch [Pag13]), albeit in a different way. Our work is the first to surpass the linear $1/r$ speed-accuracy tradeoff (AMM error) for AMM, using nothing but FFTs.
One suboptimal aspect of Theorem 1.3 is that it is agnostic to the structure of the polynomials being convolved (the construction used in Theorem 1.2). To address this, we suggest designing low degree polynomials which satisfy an approximate version of the triple product property [CU03], that captures when convolutions can embed matrix multiplication. In the abelian case, the approximate notion is equivalent to finding sumsets which minimize the (multi)-intersection with arithmetic progressions. We prove a lower bound on this problem in the abelian case, and complement it with a simple tight construction in the group $\mathbb{Z}_{rn^{2}}$. In the non-abelian case, we extend a proof of Gowers to obtain a weaker lower bound.
We propose a natural family of randomized algorithms that achieve state-of-the-art [^2] Up to a log factor in runtime. approximation (linear speed-accuracy tradeoff), using a single convolution in a small abelian group to compute an approximation of a single matrix product (as opposed to Theorem 1.3, where each convolution encodes multiple matrix products, using the rich combinatorial structure of multi-variate polynomials).
@@
Theorem 1.4 generalizes the $\Omega(n^{3})$ lower bound of [CU03] on the size of abelian groups realizing exact matrix multiplication using convolutions, and provides an alternative, elementary algorithm achieving (AMM error), while possibly having better qualities such as high output rank. Note that this result does not rule out gaining improvements by computing multiple products in a single convolution (akin to Theorem 1.3). The upside of this approach is that it is non-asymptotic at all.
Finally, we suggest a generalization of this notion to the simultaneous TPP setup, which captures a convolution’s capacity to encode multiple independent matrix multiplications via multi-variate polynomials [Coh+05]. We propose studying low-degree approximations of multi-variate polynomials, in hope of breaking the state-of-the-art while remaining non-asymptotic.
1.2 Technical Overview
@@
Matrix Multiplication via Convolutions.
@@
Unfortunately, the convolution algorithm in (1.1) is quite useless, as the polynomial $P_{\mathbf{A}}$ has degree $n^{3}$, hence the convolution (via FFT) requires $\Omega(n^{3}\log n)$ time, slower than naïve matrix multiplication(!). As such, [CU03] generalized the group-convolution approach (1.1) and the TPP property to general groups, where instead of polynomials, the matrices $\mathbf{A}$ and $\mathbf{B}$ are embedded into the “group algebra" $\mathbb{C}[G]$ of some general but fixed group $G$, and the two embeddings are then convolved inside $\mathbb{C}[G]$ to produce an element of $\mathbb{C}[G]$ containing the desired output $\mathbf{A}\mathbf{B}$. [^3] More formally, embed the indices $\mathbf{A}(i,j),\mathbf{B}(j^{\prime},k)\to G$ and compute $(\sum_{g\in G}\mathbf{A}_{g}g)\cdot(\sum_{h\in G}\mathbf{B}_{h}h)=\sum_{f\in G}(\sum_{gh=f}\mathbf{A}_{g}\mathbf{B}_{h})f$. The TPP condition for general groups then becomes:
@@
The aforementioned $\Omega(n^{3})$ lower bound on the degree of TPP polynomials in (1.1) is not a coincidence – A simple proof [CU03] shows that any abelian group satisfying the TPP must have $\left|G\right|\geq n^{3}$, which seems to make abelian groups useless.
@@
The TPP lower bound for abelian groups shifted the focus of the group-theoretic-FMM line of work to finding non-abelian TPP groups of sub-cubic size [^4] In fact, an example of a group $G$ satisfying the TPP with $\left|G\right|=n^{2+o(1)}$ was given in [CU03]. [CU03, Coh+05]. Alas, non-abelian groups generally lack fast convolution algorithms, due to: (i) the absence of a fast Fourier transform algorithm; (ii) The fact that Fourier transform of non-abelian groups transforms convolutions into block-diagonal matrix products, as opposed to element-wise product in the abelian case. This is where divide-and-conquer comes into play: applying the algorithm recursively on tiles (a-la Strassen) amortizes the cost of any linear transformation done before or after the recursive step, so applying the Fourier transform of $G$ is asymptotically free of charge [^5] This was explained in a previous footnote, discussing the computation of $\mathbf{U}^{\otimes k}v$.. This is also where our approach departs from the group-theoretic line of work – Since we aim for non-recursive (“combinatorial”) algorithms, we must inevitably work with abelian groups which have FFTs. While this seems like a dead-end in light of the $\Omega(n^{3})$ degree lower bound for TPP polynomials, it turns out that multivariate polynomials (e.g., convolutions in $\mathbb{Z}_{m}^{k}$) satisfy a more general TPP condition, proposed in [Coh+05], which allows to multiply independent block-matrices, and thus circumvents the $\Omega(n^{3})$ barrier. The property is defined by:
Definition 1.6 (STPP).
We say that a finite group $G$ satisfies the Simultaneous TPP for $\left\langle{n_{1},m_{1},p_{1}}\right\rangle,\ldots\left\langle{n_{\ell},m_{\ell},p_{\ell}}\right\rangle$, if there exist triplets of subsets in $G$, $\left\{(S_{i},T_{i},U_{i})\right\}_{i=1,\ldots,\ell}$, of sizes $\left|S_{i}\right|=n_{i},\left|T_{i}\right|=m_{i},\left|U_{i}\right|=p_{i}$, such that $(S_{i},T_{i},U_{i})$ satisfy the TPP, and they don’t mix:
Simply stated, the STPP is met when $G$ contains $\ell$ triplets $(S_{i},T_{i},U_{i})$ which marginally satisfy the (TPP). and which don’t mix together under the convolution of $G$.
1.2.1 Exact Matrix Multiplication (Proof of Theorem 1.2)
A simple construction of [Coh+05] gives a non-trivial abelian group that satisfies the (STPP).. In particular, they showed that the group $\mathbb{Z}_{m}^{3N}$ ($3N$-variate polynomials with degree at smaller than $m$ on each variable) contains $\ell=2^{N}$ triples of size $q=(m-1)^{N}$, allowing to compute $2^{N}$ independent matrix products of size $(m-1)^{N}\times(m-1)^{N}$ in a single convolution. In its most basic form, by setting $N=1$ (so d=2,q=(m1)d=2,q=(m-1)), it allows embedding two independent matrix products $\mathbf{A}_{1}\times\mathbf{B}_{1}$ and $\mathbf{A}_{2}\times\mathbf{B}_{2}$, of size $m-1\times m-1$ via the convolution of the following polynomials in $\mathbb{Z}_{m}^{3}$: